OpenStudy (anonymous):

Prove: tanx=(sinx+sin^2x)/(cosx(1+sinx)) The steps to get the answer please :)

Parth (parthkohli):

\[\dfrac{\sin(x) + \sin^2(x)}{\cos(x) (1 + \sin(x))} = \dfrac{\sin(x)(1 + \sin(x))}{\cos(x) (1 + \sin(x))}\]

Parth (parthkohli):

Can you do the rest? Hint:\[\tan(x) = \dfrac{\sin(x)}{\cos(x)}\]

Parth (parthkohli):

@isuckatmath9999

OpenStudy (anonymous):

I'm asking for the answer, not for a prompt.

OpenStudy (anonymous):

but im sure i'll figure it out